# AI Chatbots as Professional Service Agents: Developing a Professional Identity

### v1

@@ Abstract
With the rapid expansion of large language model (LLM) applications, there is an emerging shift in the role of LLM-based AI chatbots from serving merely as general inquiry tools to acting as professional service agents. However, current studies often overlook a critical aspect of professional service agents: the act of communicating in a manner consistent with their professional identities. This is of particular importance in the healthcare sector, where effective communication with patients is essential for achieving professional goals, such as promoting patient well-being by encouraging healthy behaviors. To bridge this gap, we propose LAPI (LLM-based Agent with a Professional Identity), a novel framework for designing professional service agent tailored for medical question-and-answer (Q&A) services, ensuring alignment with a specific professional identity. Our method includes a theory-guided task planning process that decomposes complex professional tasks into manageable subtasks aligned with professional objectives and a pragmatic entropy method designed to generate professional and ethical responses with low uncertainty. Experiments on various LLMs show that the proposed approach outperforms baseline methods, including few-shot prompting, chain-of-thought prompting, across key metrics such as fluency, naturalness, empathy, patient-centricity, and ROUGE-L scores. Additionally, the ablation study underscores the contribution of each component to the overall effectiveness of the approach.
@@ 1 Introduction
AI chatbots have gained great popularity worldwide to serve as professional assistants and provide services across various sectors Dam et al. (2024); Xi et al. (2023). Integrating AI chatbots is a cost-effective strategy for organizations, delivering value to customers while reducing operational expenses. For instance, automation in customer interactions is projected to save retailers approximately \$439 billion annually by 2023, as they replace many human-operated customer service roles with chatbots Williams (2019). The emergent of generative AI, particularly large language models (LLMs), is revolutionizing the functionality of chatbots by significantly enhancing their interactive communication capabilities. LLMs, such as OpenAI’s GPT-series and Llama, enabling chatbots to understand and generate natural language with remarkable fluency and context-awareness  Achiam et al. (2023); Touvron et al. (2023a).
@@
Recently, there is a trend of a significant transition in chatbot applications, shifting from general inquiry tools to professional service agents capable of delivering specialized support. The increasingly prevalent human-AI interactions in high-stake industries (e.g., healthcare and finance) require professional service agents to offer a more professional and functional experience than simple knowledge-enhanced inquiry tools Safi et al. (2020). This leads to a critical challenge in human-AI communication that has not been well addressed in existing studies, namely how AI chatbots can communicate with humans and answer questions in a manner consistent with the chatbots’ professional identity. Beyond simply responding to user queries, professional service agents always have a defined professional identity, such as that of a virtual doctor, which requires their responses to be aligned with professional goals, such as supporting patients’ health behavior change (e.g., treatment or medication adherence and reducing substance abuse)Clark and Bailey (2024). However, existing AI chatbots primarily emphasize information retrieval and question answering, overlooking the professional intention and goals that are integral to professional fields  Lee et al. (2024); Singhal et al. (2025).
@@
To develop a professional identity, an AI chatbot should have the ability to think and act like a professional. Firstly, the ability to analyze and break down complex professional tasks into more manageable components is essential. In contrast to mathematical and logical problems, which typically have clear chains of thought and definitive answers Wei et al. (2022), many question-and-answer (Q&A) tasks in a service scenario are open-ended and require free-form responses Megahed et al. (2024); Singhal et al. (2025). The challenge lies in effectively guiding these chatbots to adopt a professional thought process. Secondly, given their role as frontline professionals, AI chatbots must have the competence to effectively address user inquiries and deliver reliable services. However, it should be noted that LLMs are non-deterministic, meaning that the same input can generate different outputs, some of which may be incorrect or hallucinated Farquhar et al. (2024). These drawbacks can result in occasional inaccurate answers or hallucination problems, which is fatal in the high-stakes field Ji et al. (2023). Interactions with such chatbots have the potential to trigger undesirable behaviors in users, such as noncompliance among patients, which can negatively impact both service providers and users.
@@
Motivated by to urgent need to provide professional service to users through LLM-based chatbots, our goal is to propose a novel framework, i.e., LAPI (LLM-based Agent with a Professional Identity), which aims to enhance the effectiveness of LLM-based chatbots in serving as professional service agents, particularly in the context of medical Q&A services. To address the aforementioned challenges, we consider a subfield of linguistics, namely pragmatics, which how language is used for social interactions and the attainment of goals Ciccia and Turkstra (2002) and the use of information from context to determine the effective means of achieving a given goal Grice (1975). From a pragmatic perspective, AI chatbots should provide useful responses and remove meaningless or incorrect information. The proposed approach involves two key steps: first, the comprehension and decomposition of complex professional tasks into manageable subtasks through a theory-guided task planning process; and second, the iterative updating and eventual generation of professional and ethical responses using a pragmatic entropy method. Specifically, we propose a theory-guided task planning process that employs a conceptual framework in the healthcare domain—namely, the Health Belief Model (HBM)—to guide AI chatbots in how they think and act to support interventions aimed at encouraging healthy behaviors. Furthermore, we propose a pragmatic entropy method that generates professional and ethical responses, with the objective of maximizing the clarity, empathy, and relevance of the chatbot’s outputs while ensuring alignment with ethical standards and professional goals.
@@ 2.1 Medical AI Chatbots
The medical field represents a domain in which the implementation of AI chatbots is gaining traction as a means of facilitating access to information from the patient perspective and alleviating the workload on doctors Safi et al. (2020). A variety of methods and techniques have been employed in the development of medical chatbots for diverse objectives, including statistical methods, deep learning, and LLMs Mihailidis et al. (2008); Jothi et al. (2022); Lee et al. (2024). These bots have been developed for diverse applications, including various aspects of health and medicine. These range from chatbots that support general health and activities of daily living to chatbots that assist specific disease management, such as heart disease and cancerWong et al. (2011); Lee et al. (2024). Prior to the advent of LLMs, conventional chatbots exhibited constrained contextual and language understanding, resulting in inaccurate responses and suboptimal interaction in a human-like manner. Recently, significant advancements have been made in the development of LLM-based medical chatbots, which are designed to facilitate human-like interactions with patients Singhal et al. (2025).
@@ 2.2 Professional LLM-based Chatbots
Chatbots have stepped into the age of LLM empowerment. The rapid development of LLMs enhances the capabilities of chatbots by improving their conversational proficiency and facilitating human-like interactions. ChatGPT exemplifies this category of chatbot, optimized to generate natural, human-like dialogue. Other widely used LLMs include GPT4, Llama, Mistral, ChatGLM2, and Gemma Achiam et al. (2023); Touvron et al. (2023a); Jiang et al. (2023); Zeng et al. (2024); Mesnard et al. (2024). Chatbots have been employed in customer service to address customer inquiries regarding products or services, either as an initial point of contact or as an alternative to speaking with a human representative Hong et al. (2024). In addition to healthcare, LLM-based chatbots have gradually become powerful tools in areas such as retail, education, research, and many others Megahed et al. (2024); Benzinho et al. (2024); Giudici et al. (2024). To improve chatbot’s ability on professional tasks, such as programming education and home designing, various knowledge-enhanced strategies, such as fine-tuning, prompt engineering, and Retrieval-Augmented Generation (RAG), have been employed Lewis et al. (2020).
@@ 3.1 Professional Service Agents
From a philosophical standpoint, the term ”agent” traditionally describes entities possessing desires, beliefs, intentions, and the capacity to action Xi et al. (2023); Davidson (1963). Human agents, for instance, are frequently employed by companies to handle customer service operations, focusing on addressing customer queries and ensuring satisfaction through effective communication and problem-solving skills. With the progression of AI, the concept of an ”agent” has expanded to describe AI entities that are characterized by intelligent behavior, such as autonomy, proactiveness, and social capabilities. AI agents, particularly chatbots powered by large language models (LLMs), have demonstrated remarkable versatility across various scenarios, attributable to their extensive inherent knowledge and human-like communication abilities Huang et al. (2024); Chong et al. (2021).
LLM-based agents are typically structured around three key components. Agent Core serves as the main brain of the system, typically utilizing an LLM with general-purpose capabilities. Task Planning is responsible for decomposing complex tasks into manageable subtasks, thereby enhancing the agent’s reasoning and reliability of responses. Actions of a chatbot are communication with humans. A professional AI agent’s primary actions are aimed at high-level information transmission to fulfill a robot’s intent, a process referred to as “intention action.” A typical example of such agents is applications in the healthcare domain, where empathy-aware agents assist in diagnostics and knowledge retrieval to provide a user-friendly service Mao et al. (2022).
@@
Responding to the trend of chatbots shifting from general inquire tools to professional service agents, we identify the significant distinction between AI-chatbots as inquiry tools and as professional service agents, as shown in Figure 1. Professional identity is a crucial attribute of a professional service agent, significantly shaping the agent’s goals, task planning, and action execution. Chatbots with professional identity have distinct professional intentions and goals, which in turn influence their task planning and action execution. Using medical Q&A as an example, inquiry tools primarily focus on answering medical questions Safi et al. (2020). In contrast, professional service agents serve two purposes: responding to medical inquiries and promoting patient well-being by encouraging healthy behaviors, aligning with their professional identity (e.g., doctor). For task planning, the decomposition of complex professional tasks should adhere to established professional standards, informed by either theories or expert guidelines. Regarding actions, the chatbot’s responses should align with professional goals. This alignment ensures that the chatbot not only addresses patient questions but also offers suggestions and persuasive guidance to foster healthy behaviors.
@@ 3.2 Health Belief Model
The Health Belief Model (HBM) is a widely used conceptual framework in health behavior research, which supports interventions aimed at encouraging healthy behaviorsChampion and Skinner (2008). Initially proposed in the 1950s, the model has evolved in response to practical public health concern and healthcare management. Rooted in a well-established body of psychological and behavioral theories, the HBM posits that health-related behavior depends primarily influenced by two factors: (1) the individual’s desire to avoid illness or recover from it if already affected; and (2) the belief that a particular health action care prevent or mitigate the illness Janz and Becker (1984).
@@
The HBM has been employed to inform health intervention designs aimed at promoting healthy behaviors Orji et al. (2012), particularly as patient-centric (proactive) healthcare gains popularity. Previous research has shown that interventions grounded in established theories and models tend to be more effective than those developed based on intuition Glanz (1997).
@@ 4.1 Problem Formulation
Given a patient’s question $q$ and a set of predefined pragmatic rules aligning with a professional identity $R=\{r_{1},r_{2},r_{3},\dots,r_{m}\}$, the objective is to generate a response $s$ that not only provides an professional answer to the query but also adheres to these established rules, thereby reflecting the professional identity (i.e., the doctor in our experiment setting). The desired response $s^{*}$ should satisfy the following condition: it adheres to the pragmatic rules with a minimal level of uncertainty in the response distribution. The problem can be formally expressed as:
$s^{*}=\text{LLM}(q,R,\tau,\epsilon)$ (1)
where $\tau$ is the satisfaction rate threshold, and $\epsilon$ is the pragmatic entropy threshold. The final response $s^{*}$ is selected when the satisfaction rate $P(C_{\text{all}}|S)\geq\tau$ and the pragmatic entropy $PE(S)\leq\epsilon$, ensuring both accuracy and alignment with the predefined professional characteristics.
@@ 4.2 Theory-guided Task Planning
Based on the prompt shown in Figure 3, the LLM assesses the query’s relevance to each category and assigns a corresponding weight to each. Next, we generate the initial prompt $p_{\text{initial}}$ for the LLM by constructing two specific prompts: one fully focused on the Perceived Threat category, and another fully focused on Perceived Effectiveness of Health Behavior. These prompts are pre-configured to ensure that the LLM generates an initial prompt that accurately reflects the correct placement of the question within the HBM theory framework. The content of these prompts is in Appendix A.
@@ 4.3 Pragmatic Entropy
Pragmatic entropy aims to assist the AI chatbot generate professional and ethical response by 1) iteratively optimizing a given prompt $p$ to generate responses that maximize the proportion of outputs satisfying all predefined pragmatic rules $R=\{r_{1},r_{2},r_{3},...,r_{m}\}$, and 2) selecting responses with minimized uncertainty from the category that meet all pragmatic rules. Pragmatic entropy measures the variability in the distribution of responses across different rule-satisfaction categories, with lower entropy indicating more focused outputs that align with the desired rules. By minimizing entropy, the method ensures stable and reliable generation, improving alignment with professional goals.
Response Evaluation.
Given a prompt $p$, the LLM generates a set of $n$ responses by sampling from its output distribution with temperature 1, which ensures a balance between randomness and determinism in the generated outputs:
@@
where $S=\{s_{1},s_{2},\dots,s_{n}\}$ represents the generated responses. The temperature parameter controls the randomness of the sampling process. A temperature of $1$ maintains the default probability distribution of the model, allowing diverse but coherent outputs.
Each response $s\in S$ is classified into categories $C_{j}$ based on its satisfaction of specific subsets of the pragmatic rules $R$. This classification process is judged by the Large Language Model (LLM) itself. The following metrics are computed to evaluate the quality of the prompt:
  1. 1.
    Satisfaction Rate: The proportion of responses satisfying all rules is given by:
@@
    where |$\left|\{s_{i}\in S\mid s_{i}\in C_{\text{all}}\}\right|$ denotes the number of responses in $S$ that fall into category $C_{\text{all}}$, and $n$ is the total number of generated responses.
  2. 2.
    Pragmatic Entropy: The entropy of the output distribution across all categories $C_{j}$ is computed as:
@@
    where $P(C_{j}|S)=\frac{\left|\{s_{i}\in S\mid s_{i}\in C_{j}\}\right|}{n}$ represents the proportion of responses in $S$ that fall into category $C_{j}$.
Optimization Process
@@
To improve the prompt, the current prompt $p$ along with information on unmet conditions is fed into the large language model (LLM) to generate a refined version $p_{\text{new}}$:
$p_{\text{new}}=\text{LLM}(p,unmet\_conditions,\text{temperature}=0.3)$ (5)
where $p_{\text{new}}$ is the updated prompt. Unmet conditions refer to the missing pragmatic rules identified within the category that has the largest proportion in categories that have not satisfied all the rules. The parameter $\text{temperature}=0.3$ Moslem et al. (2023); Wassie (2024) is chosen to balance stability and flexibility in the prompt generation process.
The refined prompt $p_{\text{new}}$ is evaluated quantitatively by generating a new set of $n$ responses:
$S_{\text{new}}=\text{LLM}(p_{\text{new}},\text{temperature}=1).$ (6)
Each response $s\in S_{\text{new}}$ is classified into categories $C_{j}$, and the satisfaction rate $P(C_{\text{all}}|S_{\text{new}})$ and pragmatic entropy $PE(S_{\text{new}})$ are recomputed.
If the satisfaction rate $P(C_{\text{all}}|S)$ meets or exceeds a predefined threshold $\tau$, and the pragmatic entropy $PE(S)$ is below a threshold $\epsilon$, the prompt $p$ is considered optimal, and the process terminates with $p^{*}=p$. Otherwise, the model proceeds to refine the prompt.
Algorithm 1 Pragmatic entropy optimization process
@@
  Output: Final response answer.
@@
4:     Compute satisfaction rate $P(C_{\text{all}}|S)$ using Eq.(3);
5:     Compute pragmatic entropy $PE(S)$ based on Eq.(4);
6:     Select $s^{*}$ randomly from $S$;
@@
10:        Identify $unmet\_conditions$ from categories that have not satisfied all the rules;
11:        Get the updated prompt $p_{\text{new}}$ using Eq.(5);

### v2

@@ Abstract
With the rapid expansion of large language model (LLM) applications, there is an emerging shift in the role of LLM-based AI chatbots from serving as general inquiry tools to acting as professional service agents. However, current studies often overlook a critical aspect of professional service agents: the act of communicating in a manner consistent with their professional identities. This is of particular importance in the healthcare sector, where effective communication with patients is essential for achieving professional goals, such as promoting patient well-being by encouraging healthy behaviors. To bridge this gap, we propose LAPI (LLM-based Agent with a Professional Identity), a novel framework for designing professional service agent tailored for medical question-and-answer (Q&A) services, ensuring alignment with a specific professional identity. Our method includes a theory-guided task planning process that decomposes complex professional tasks into manageable subtasks aligned with professional objectives and a pragmatic entropy method designed to generate professional and ethical responses with low uncertainty. Experiments on various LLMs show that the proposed approach outperforms baseline methods, including few-shot prompting, chain-of-thought prompting, across key metrics such as fluency, naturalness, empathy, patient-centricity, and ROUGE-L scores. Additionally, the ablation study underscores the contribution of each component to the overall effectiveness of the approach.
@@ 1 Introduction
AI chatbots have gained great popularity worldwide to serve as professional assistants and provide services across various sectors[dam-etal:llm-chatbots, xi-etal:llm-agents]. Integrating AI chatbots is a cost-effective strategy for organizations, delivering value to customers while reducing operational expenses. For instance, automation in customer interactions is projected to save retailers approximately \$439 billion annually by 2023, as they replace many human-operated customer service roles with chatbots [williams:chatbots-retail-sales]. The emergent of generative AI, particularly large language models (LLMs), is revolutionizing the functionality of chatbots by significantly enhancing their interactive communication capabilities. LLMs, such as OpenAI’s GPT-series and Llama, enabling chatbots to understand and generate natural language with remarkable fluency and context-awareness [achiam-etal:gpt-4-technical-report, touvron-etal:llama-foundation-models].
@@
Recently, there is a trend of a significant transition in chatbot applications, shifting from general inquiry tools to professional service agents capable of delivering specialized support. The increasingly prevalent human-AI interactions in high-stake industries (e.g., healthcare and finance) require professional service agents to offer a more professional and functional experience than simple knowledge-enhanced inquiry tools[safi-etal:chatbots-medical-applications]. This leads to a critical challenge in human-AI communication that has not been well addressed in existing studies, namely how AI chatbots can communicate with humans and answer questions in a manner consistent with the chatbots’ professional identity. Beyond simply responding to user queries, professional service agents always have a defined professional identity, such as that of a virtual doctor, which requires their responses to be aligned with professional goals, such as supporting patients’ health behavior change (e.g., treatment or medication adherence and reducing substance abuse)[clark-bailey:chatbots-healthcare]. However, existing AI chatbots primarily emphasize information retrieval and question answering, overlooking the professional intention and goals that are integral to professional fields [lee-etal:ai-generated-medical-responses, singhal-etal:expert-level-medical-question-answering].
@@
To develop a professional identity, an AI chatbot should have the ability to think and act like a professional. Firstly, the ability to analyze and break down complex professional tasks into more manageable components is essential. In contrast to mathematical and logical problems, which typically have clear chains of thought and definitive answers[wei-etal:chain-of-thought], many question-and-answer (Q&A) tasks in a service scenario are open-ended and require free-form responses[megahed-etal:misuse-of-ai-in-spc, singhal-etal:expert-level-medical-question-answering]. The challenge lies in effectively guiding these chatbots to adopt a professional thought process. Secondly, given their role as frontline professionals, AI chatbots must have the competence to effectively address user inquiries and deliver reliable services. However, it should be noted that LLMs are non-deterministic, meaning that the same input can generate different outputs, some of which may be incorrect or hallucinated[farquhar-etal:detecting-hallucinations]. These drawbacks can result in occasional inaccurate answers or hallucination problems, which is fatal in the high-stakes field[ji-etal:survey-hallucination-nlg]. Interactions with such chatbots have the potential to trigger undesirable behaviors in users, such as noncompliance among patients, which can negatively impact both service providers and users.
@@
Motivated by to urgent need to provide professional service to users through LLM-based chatbots, our goal is to propose a novel framework, i.e., LAPI (LLM-based Agent with a Professional Identity), which aims to enhance the effectiveness of LLM-based chatbots in serving as professional service agents, particularly in the context of medical Q&A services. To address the aforementioned challenges, we consider a subfield of linguistics, namely pragmatics, which how language is used for social interactions and the attainment of goals[ciccia:cohesion-communication-burden] and the use of information from context to determine the effective means of achieving a given goal[grice:logic-conversation]. From a pragmatic perspective, AI chatbots should provide useful responses and remove meaningless or incorrect information. The proposed approach involves two key steps: first, the comprehension and decomposition of complex professional tasks into manageable subtasks through a theory-guided task planning process; and second, the iterative updating and eventual generation of professional and ethical responses using a pragmatic entropy method. Specifically, we propose a theory-guided task planning process that employs a conceptual framework in the healthcare domain—namely, the Health Belief Model (HBM)—to guide AI chatbots in how they think and act to support interventions aimed at encouraging healthy behaviors. Furthermore, we propose a pragmatic entropy method that generates professional and ethical responses, with the objective of maximizing the clarity, empathy, and relevance of the chatbot’s outputs while ensuring alignment with ethical standards and professional goals.
@@ 2.1 Medical AI Chatbots
The medical field represents a domain in which the implementation of AI chatbots is gaining traction as a means of facilitating access to information from the patient perspective and alleviating the workload on doctors[safi-etal:chatbots-medical-applications]. A variety of methods and techniques have been employed in the development of medical chatbots for diverse objectives, including statistical methods, deep learning, and LLMs[mihailidis-etal:coach-prompting, jothi-etal:humanoid-chatbot-medical, lee-etal:ai-generated-medical-responses]. These bots have been developed for diverse applications, including various aspects of health and medicine. These range from chatbots that support general health and activities of daily living to chatbots that assist specific disease management, such as heart disease and cancer[wong-etal:health-conversational-system, lee-etal:ai-generated-medical-responses]. Prior to the advent of LLMs, conventional chatbots exhibited constrained contextual and language understanding, resulting in inaccurate responses and suboptimal interaction in a human-like manner. Recently, significant advancements have been made in the development of LLM-based medical chatbots, which are designed to facilitate human-like interactions with patients[singhal-etal:expert-level-medical-question-answering].
@@ 2.2 Professional LLM-based Chatbots
Chatbots have stepped into the age of LLM empowerment. The rapid development of LLMs enhances the capabilities of chatbots by improving their conversational proficiency and facilitating human-like interactions. ChatGPT exemplifies this category of chatbot, optimized to generate natural, human-like dialogue. Other widely used LLMs include GPT4, Llama, Mistral, ChatGLM2, and Gemma[achiam-etal:gpt-4-technical-report, touvron-etal:llama-foundation-models, jiang-etal:mistral-7b, glm-etal:chatglm, gemma-etal:gemma-open-models]. Chatbots have been employed in customer service to address customer inquiries regarding products or services, either as an initial point of contact or as an alternative to speaking with a human representative[hong-etal:expanding-chatbot-knowledge]. In addition to healthcare, LLM-based chatbots have gradually become powerful tools in areas such as retail, education, research, and many others[megahed-etal:misuse-of-ai-in-spc, benzinho-etal:llm-chatbot-farm-to-fork, giudici-etal:home-automation-llm-chatbot]. To improve chatbot’s ability on professional tasks, such as programming education and home designing, various knowledge-enhanced strategies, such as fine-tuning, prompt engineering, and Retrieval-Augmented Generation (RAG), have been employed[lewis-etal:retrieval-augmented-generation].
@@ 3.1 Professional Service Agents
From a philosophical standpoint, the term "agent" traditionally describes entities possessing desires, beliefs, intentions, and the capacity to action[xi-etal:llm-agents, donald:actions-reasons-causes]. Human agents, for instance, are frequently employed by companies to handle customer service operations, focusing on addressing customer queries and ensuring satisfaction through effective communication and problem-solving skills. With the progression of AI, the concept of an "agent" has expanded to describe AI entities that are characterized by intelligent behavior, such as autonomy, proactiveness, and social capabilities. AI agents, particularly chatbots powered by large language models, have demonstrated remarkable versatility across various scenarios, attributable to their extensive inherent knowledge and human-like communication abilities[huang-etal:agent-ai-holistic-intelligence, chong-etal:ai-chatbots-services-frontline].
LLM-based agents are typically structured around three key components. Agent Core serves as the main brain of the system, typically utilizing an LLM with general-purpose capabilities. Task Planning is responsible for decomposing complex tasks into manageable subtasks, thereby enhancing the agent’s reasoning and reliability of responses. Actions of a chatbot are communication with humans. A professional AI agent’s primary actions are aimed at high-level information transmission to fulfill a robot’s intent, a process referred to as “intention action.” A typical example of such agents is applications in the healthcare domain, where empathy-aware agents assist in diagnostics and knowledge retrieval to provide a user-friendly service[mao-etal:biases-of-pretrained-models].
@@
Responding to the trend of chatbots shifting from general inquire tools to professional service agents, we identify the significant distinction between AI-chatbots as inquiry tools and as professional service agents, as shown in Figure 1. Professional identity is a crucial attribute of a professional service agent, significantly shaping the agent’s goals, task planning, and action execution. Chatbots with professional identity have distinct professional intentions and goals, which in turn influence their task planning and action execution. Using medical Q&A as an example, inquiry tools primarily focus on answering medical questions[safi-etal:chatbots-medical-applications]. In contrast, professional service agents serve two purposes: responding to medical inquiries and promoting patient well-being by encouraging healthy behaviors, aligning with their professional identity (e.g., doctor). For task planning, the decomposition of complex professional tasks should adhere to established professional standards, informed by either theories or expert guidelines. Regarding actions, the chatbot’s responses should align with professional goals. This alignment ensures that the chatbot not only addresses patient questions but also offers suggestions and persuasive guidance to foster healthy behaviors.
@@ 3.2 Health Belief Model
The Health Belief Model (HBM) is a widely used conceptual framework in health behavior research, which supports interventions aimed at encouraging healthy behaviors[champion-skinner:health-belief-model]. Initially proposed in the 1950s, the model has evolved in response to practical public health concern and healthcare management. Rooted in a well-established body of psychological and behavioral theories, the HBM posits that health-related behavior depends primarily influenced by two factors: (1) the individual’s desire to avoid illness or recover from it if already affected; and (2) the belief that a particular health action care prevent or mitigate the illness[janz-becker:health-belief-model].
@@
The HBM has been employed to inform health intervention designs aimed at promoting healthy behaviors[orji-etal:health-interventions-health-belief-model], particularly as patient-centric (proactive) healthcare gains popularity. Previous research has shown that interventions grounded in established theories and models tend to be more effective than those developed based on intuition[glanz:theory-at-a-glance].
@@ 4.1 Problem Formulation
Given a patient’s question $q$ and a set of predefined pragmatic rules aligning with a professional identity $R=\{r_{1},r_{2},r_{3},\dots,r_{m}\}$, the objective is to generate a response $s$ that not only provides an professional answer to the query but also adheres to these established rules, thereby reflecting the professional identity (i.e., the doctor in our experiment setting). The desired response $s$ should satisfy the following condition: it adheres to the pragmatic rules with a minimal level of uncertainty in the response distribution. The problem can be formally expressed as:
$s=\text{LLM}(q,R,\tau,\epsilon)$ (1)
where $\tau$ is the satisfaction rate threshold, and $\epsilon$ is the pragmatic entropy threshold.
@@ 4.2 Theory-guided Task Planning
Based on the prompt shown in Figure 4 of Appendix A, the LLM assesses the query’s relevance to each category and assigns a corresponding weight to each. Next, we generate the initial prompt $p_{\text{initial}}$ for the LLM by constructing two specific prompts: one fully focused on the Perceived Threat category, and another fully focused on Perceived Effectiveness of Health Behavior. These prompts are pre-configured to ensure that the LLM generates an initial prompt that accurately reflects the correct placement of the question within the HBM theory framework. The content of the prompt is detailed in Figure 5 of Appendix A.
@@ 4.3 Pragmatic Entropy
Pragmatic entropy aims to assist the AI chatbot in generating professional and ethical responses by reducing the entropy of model outputs across different pragmatic rule categories and maximizing the proportion of responses that satisfy all predefined pragmatic rules $R=\{r_{1},r_{2},r_{3},...,r_{m}\}$.
Given a prompt $p$, the LLM generates a set of $n$ responses by sampling from its output distribution with temperature 1, which enables the model’s outputs to more authentically reflect its internal representations and reasoning processes:
@@
where $S=\{s_{1},s_{2},\dots,s_{n}\}$ represents the generated responses.
Each response $s\in S$ is classified into categories $C_{j}$ based on its satisfaction of specific subsets of the pragmatic rules $R$. This classification process is judged by the LLM itself. The following metrics are computed to evaluate the quality of the prompt.
The proportion of responses satisfying all rules is given by:
@@
where $\left|\{s_{i}\in S\mid s_{i}\in C_{\text{all}}\}\right|$ denotes the number of responses in $S$ that fall into category $C_{\text{all}}$, and $n$ is the total number of generated responses.
The entropy of the output distribution across all categories $C_{j}$:
@@
where $P(C_{j}|S)=\frac{\left|\{s_{i}\in S\mid s_{i}\in C_{j}\}\right|}{n}$ represents the proportion of responses in $S$ that fall into $C_{j}$.
@@
To improve the prompt, the current prompt $p$ along with information on unmet conditions (Ucs) is fed into the large language model to generate a refined version $p_{\text{new}}$:
$p_{\text{new}}=\text{LLM}(p,Ucs,\text{temperature}=0.3)$ (5)
where $p_{\text{new}}$ is the updated prompt. Ucs refer to the missing pragmatic rules identified within the category that has the largest proportion in categories that have not satisfied all the rules. The parameter $\text{temperature}=0.3$[moslem2023finetuninglargelanguagemodels, wassie2024machinetranslationgeezlanguage] is chosen to balance stability and flexibility in the prompt generation process.
If the satisfaction rate $P(C_{\text{all}}|S)$ meets or exceeds a predefined threshold $\tau$, and the pragmatic entropy $PE(S)$ is below a threshold $\epsilon$, the prompt $p$ is considered optimal, and the process terminates with $p^{*}=p$. Otherwise, the model proceeds to refine the prompt. The overall process can be seen in Algorithm 1.
Algorithm 1 Pragmatic entropy
@@
Output: Final response $s^{*}$.
@@
4:  Compute $P(C_{\text{all}}|S)$ using Eq.(3);
5:  Compute $PE(S)$ based on Eq.(4);
6:  Select $s^{*}$ randomly from $C_{\text{all}}$;
@@
10:   Identify UcsUcs from categories that have not satisfied all the rules;
11:   Get prompt $p_{\text{new}}$ using Eq.(5);